OpenStudy (idealist10):

An object stretches a spring 5 cm in equilibrium. It is initially displaced 10 cm above equilibrium and given an upward velocity of 0.25 m/s. Find its displacement for t>0. Find the frequency, period, amplitude, and phase angle of the motion.

OpenStudy (idealist10):

\[y''+\frac{ k }{ m }y=0\]

OpenStudy (idealist10):

How do I find k/m?

OpenStudy (idealist10):

@freckles @Kainui

OpenStudy (irishboy123):

from the initial equilibrium displacement, you have \[mg = kx_o\] \[\dfrac{k}{m} = \dfrac{g}{x_o}\] does that get the ball rolling?